7. You are tasked with determining the Kow for a new compound synthesized by a pesticide company. The molecular weight for the compound is 223 g/mol. During an extraction in octanol/water you determine that there are 1.78 g of the sample in a 250 mL oil and 1.69 mg in 450 mL of water. a. What is the Kow of the compound? b. Will this compound persist in the environment? Why or why not? c. Will this compound be safe in the environment?
Answer #1

Solution:-

A) Kow :- It is the ratio of the concentration of the compound in octanol and water.

Kow = Concentration in octanol/concentration in water

so first we will have to determine the concentration of the compound in water and octanol respectively.

concentration in water: Molarity = mole/ volume in liters [ mole = weight / molar mass ]

the weight of the compound in water = 1.69 mg = 0.00169 g [ 1g = 1/1000 g ]

mole of compound in water = 0.00169g / 223 g/mol = 7.56 * 10-6 mole [molar mass of compound = 223 g/mol]

molarity in water = (( 7.56 * 10-6) * 1000 ) / 450 = 1.68 * 10-5 M [ 1 ml = 1/1000 L]

now we will calculate concentration in octanol

Concentration in octanol :

mole in octanol = 1.78 g /223 g/mol = 0.00799 mol

concentration ( Molarity ) = 0.00799 mol * 1000 / 250 = 0.032 M

Kow = concentration in octanol/ concentration in water

= 0.032 / (1.68 * 10-5 ) = 1902.4

Kow = 1902.4

B) log (Kow ) = log ( 1902.4) = 3.28

yes,the compound will persist in environment . from the concentration of the compound in water and octanol we can clearly see that the compound is very much soluble in octanol as compared to water. It is also very clear from the Kow value which is very high means compound is very less soluble in water. as the compound is very less soluble in water it will not dissolve in water and will not enter into the water system so as in the biological cycle. hence this compound will remain in the environment.

C) This compound will be safe because this is less soluble in water so it will not enter into the water body. and also the compounds with log Kow value small than 4.5 are assumed to be safe for living being and this compound has log Kow = 3.28 .